# CR-UTP: Certified Robustness against Universal Text Perturbations

### v1

@@ Abstract
CR-UTP: Certified Robustness against Universal Text Perturbations
Qian Lou1, Xin Liang1*, Jiaqi Xue1*, Yancheng Zhang1, Rui Xie2, Mengxin Zheng1 1Department of Computer Science, University of Central Florida 2Department of Statistics and Data Science, University of Central Florida
**footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Prompt-based Language Models (PLMs) Thoppilan et al. (2022); Zeng et al. (2022); Achiam et al. (2023); Touvron et al. (2023b); Chiang et al. (2023) have achieved significant success across a wide range of real-world applications Wu et al. (2020); Brown et al. (2020a); Wei et al. (2022); Chowdhery et al. (2023). However, despite their prominent performance, PLMs have been shown vulnerable to noises and perturbations on the input Xu et al. (2022); Shayegani et al. (2023); Lou et al. (2022); Al Ghanim et al. (2023); Zheng et al. (2023). Such vulnerability has notably restricted PLM’s utility, especially in high-stake environments such as bank records analysis Heaton et al. (2017), health care records analysis Myszczynska et al. (2020). In these settings, the stability of every prediction is critical, i.e., PLM predictions should remain consistent despite minor input variations, such as word substitutions Alzantot et al. (2018); Ren et al. (2019); Li et al. (2020a). This concern aligns with the study of certified robust PLMs Zeng et al. (2023), which guarantees that all PLM predictions are accurate within the local vicinity of the input.
@@ 2 Background and Motivation
Certified Robustness in Language Models. Numerous defense methods, such as adversarial training and perturbation-controlled approaches, have been developed to counteract adversarial attacks Wang et al. (2019); Zhou et al. (2021); Goyal et al. (2023). However, these traditional tools may become ineffective against novel, advanced attack strategies. To address this, certified robustness has been introduced, offering a guarantee against any attack as long as the number of perturbed words remains below a certain threshold. A model achieves certification if it can consistently produce the correct output when the number of perturbations does not exceed the certified radius. While models of smaller size can obtain robustness certification through deterministic methods Li et al. (2020b); Ostrovsky et al. (2022); Weng et al. (2018); Kolter and Wong (2017), the computational demands of language models preclude such approaches. Consequently, probabilistic methods, such as Random Smoothing Cohen et al. (2019), have been introduced to certify the robustness of large language models.
Random Smoothing. Random Smoothing, a promising approach introduced by Cohen et al. (2019); Weber et al. (2020), certifies the robustness of large neural networks. This method enhances a model’s robustness by adding Gaussian noise to the original input Salman et al. (2020); Li et al. (2023). It was quickly adopted for large language models in the NLP field, exemplified by SAFER Ye et al. (2020) and Randomized [MASK] Zeng et al. (2023). To improve model performance with random smoothing, the computer vision field has explored re-training the original model to adapt to Gaussian noise Jeong and Shin (2020); Zhai et al. (2020); Salman et al. (2019). However, applying this re-training method in NLP to achieve a model tolerant to smoothing is prohibitively expensive. Instead, in the NLP domain, Zhang et al. (2023) proposed a self de-noising method that allows the Pretrained Language Model (PLM) itself to recover the information lost due to masking.
@@ 5.2 Ablation Study
Comparison with Adversarial Training. We compare the empirical defence effects of adversarial training and our CR-UTP in Table 4. Our CR-UTP significantly reduces the ASR by over 30% while maintaining similar accuracy, outperforming adversarial training. CR-UTP consistently defends against various adversarial attacks, such as TrojLLM and TextFooler, unlike adversarial training, which shows inconsistent defense effectiveness. For instance, adversarial training effectively reduces ASR from 91.88% to 80.68% for attacks it was trained against TrojLLM, but it provides minimal defense against different attacks (TextFooler), only reducing ASR from 92.27% to 91.13%.
@@ 6 Limitation
The limitations of our paper are as follows: (i) Certified Accuracy. Although our CR-UTR has demonstrated improvements in certified accuracy and reduced ASR, achieving state-of-the-art results, there remains a gap between clean accuracy and certified accuracy. (ii) Efficiency. The CR-UTP with prompt ensemble would result in higher inference overhead compared to using just one prompt. However, it is important to note that a superior prompt can significantly enhance the effectiveness of defense strategies. Moreover, the Superior Prompt Search is an offline process to train the policy model which could be reused to generate multiple superior prompts in several seconds.

### v2

@@ Abstract
CR-UTP: Certified Robustness against Universal Text Perturbations on Large Language Models
Qian Lou1$+$, Xin Liang1*, Jiaqi Xue1*, Yancheng Zhang1, Rui Xie2, Mengxin Zheng1 1Department of Computer Science, University of Central Florida 2Department of Statistics and Data Science, University of Central Florida
++footnotetext: Corresponding author: Qian Lou, qian.lou@ucf.edu**footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Prompt-based Language Models (PLMs) Thoppilan et al. (2022); Zeng et al. (2022); Achiam et al. (2023); Touvron et al. (2023b); Chiang et al. (2023) have achieved significant success across a wide range of real-world applications Wu et al. (2020); Brown et al. (2020a); Wei et al. (2022); Chowdhery et al. (2023). However, despite their prominent performance, PLMs have been shown vulnerable to noises and perturbations on the input Xu et al. (2022); Shayegani et al. (2023); Lou et al. (2022); Al Ghanim et al. (2023); Zheng et al. (2023b). Such vulnerability has notably restricted PLM’s utility, especially in high-stake environments such as bank records analysis Heaton et al. (2017), health care records analysis Myszczynska et al. (2020). In these settings, the stability of every prediction is critical, i.e., PLM predictions should remain consistent despite minor input variations, such as word substitutions Alzantot et al. (2018); Ren et al. (2019); Li et al. (2020a). This concern aligns with the study of certified robust PLMs Zeng et al. (2023), which guarantees that all PLM predictions are accurate within the local vicinity of the input.
@@ 2 Background and Motivation
Adversarial Attacks. Adversarial attacks in deep learning involve embedding a trigger into certain training samples, thereby creating poisoned datasets. When a deep learning model is trained on such tainted datasets, it behaves normally when presented with clean inputs but exhibits malicious behavior when encountering inputs containing the trigger. In the realm of visual images, triggers often manifest as tiny patches Zheng et al. (2023c) or global perturbations Li et al. (2023a); Zheng et al. (2023a). In the context of language data, triggers can be rare words or characters like "cf" Kurita et al. (2020). This paper specifically focuses on text-based attacks.
@@
Certified Robustness in Language Models. Numerous defense methods, such as adversarial training Yoo and Qi (2021), model detection Zheng et al. (2023d) and perturbation-controlled approaches, have been developed to counteract adversarial attacks Wang et al. (2019); Zhou et al. (2021); Goyal et al. (2023). However, these traditional tools may become ineffective against novel, advanced attack strategies. To address this, certified robustness has been introduced, offering a guarantee against any attack as long as the number of perturbed words remains below a certain threshold. A model achieves certification if it can consistently produce the correct output when the number of perturbations does not exceed the certified radius. While models of smaller size can obtain robustness certification through deterministic methods Li et al. (2020b); Ostrovsky et al. (2022); Weng et al. (2018); Kolter and Wong (2017), the computational demands of language models preclude such approaches. Consequently, probabilistic methods, such as Random Smoothing Cohen et al. (2019), have been introduced to certify the robustness of large language models.
Random Smoothing. Random Smoothing, a promising approach introduced by Cohen et al. (2019); Weber et al. (2020), certifies the robustness of large neural networks. This method enhances a model’s robustness by adding Gaussian noise to the original input Salman et al. (2020); Li et al. (2023b). It was quickly adopted for large language models in the NLP field, exemplified by SAFER Ye et al. (2020) and Randomized [MASK] Zeng et al. (2023). To improve model performance with random smoothing, the computer vision field has explored re-training the original model to adapt to Gaussian noise Jeong and Shin (2020); Zhai et al. (2020); Salman et al. (2019). However, applying this re-training method in NLP to achieve a model tolerant to smoothing is prohibitively expensive. Instead, in the NLP domain, Zhang et al. (2023) proposed a self de-noising method that allows the Pretrained Language Model (PLM) itself to recover the information lost due to masking.
@@ 5.2 Ablation Study
Comparison with Adversarial Training. We compare the empirical defence effects of adversarial training Yoo and Qi (2021) and our CR-UTP in Table 4. Our CR-UTP significantly reduces the ASR by over 30% while maintaining similar accuracy, outperforming adversarial training. CR-UTP consistently defends against various adversarial attacks, such as TrojLLM and TextFooler, unlike adversarial training, which shows inconsistent defense effectiveness. For instance, adversarial training effectively reduces ASR from 91.88% to 80.68% for attacks it was trained against TrojLLM, but it provides minimal defense against different attacks (TextFooler), only reducing ASR from 92.27% to 91.13%.
@@ 6 Limitation
The limitations of our paper are as follows: (i) Certified Accuracy. Although our CR-UTR has demonstrated improvements in certified accuracy and reduced ASR, achieving state-of-the-art results, there remains a gap between clean accuracy and certified accuracy. (ii) Broader Applications. While our CR-UTP primarily concentrates on classification tasks, broadening its application to encompass other NLP tasks like generation Xue et al. (2024) would present a captivating expansion of our research. (iii) Efficiency. The CR-UTP with prompt ensemble would result in higher inference overhead compared to using just one prompt. However, it is important to note that a superior prompt can significantly enhance the effectiveness of defense strategies. Moreover, the Superior Prompt Search is an offline process to train the policy model which could be reused to generate multiple superior prompts in several seconds.